# RG-Attn: Radian Glue Attention for Multi-modality Multi-agent Cooperative Perception

### v1

@@ Abstract
Cooperative perception offers an optimal solution to overcome the perception limitations of single-agent systems by leveraging Vehicle-to-Everything (V2X) communication for data sharing and fusion across multiple agents. However, most existing approaches focus on single-modality data exchange, limiting the potential of both homogeneous and heterogeneous fusion across agents. This overlooks the opportunity to utilize multi-modality data per agent, restricting the system’s performance. In the automotive industry, manufacturers adopt diverse sensor configurations, resulting in heterogeneous combinations of sensor modalities across agents. To harness the potential of every possible data source for optimal performance, we design a robust LiDAR and camera cross-modality fusion module, Radian-Glue-Attention (RG-Attn), applicable to both intra-agent cross-modality fusion and inter-agent cross-modality fusion scenarios, owing to the convenient coordinate conversion by transformation matrix and the unified sampling/inversion mechanism. We also propose two different architectures, named Paint-To-Puzzle (PTP) and Co-Sketching-Co-Coloring (CoS-CoCo), for conducting cooperative perception. PTP aims for maximum precision performance and achieves smaller data packet size by limiting cross-agent fusion to a single instance, but requiring all participants to be equipped with LiDAR. In contrast, CoS-CoCo supports agents with any configuration—LiDAR-only, camera-only, or LiDAR-camera-both, presenting more generalization ability. Our approach achieves state-of-the-art (SOTA) performance on both real and simulated cooperative perception datasets. The code will be released at GitHub in early 2025.
@@ I INTRODUCTION
The famous phrase “United we stand, divided we fall” by Aesop aptly captures the essence of multi-agent cooperative perception. Shared and fused perception information can be the crucial stepping stone, akin to an oracle’s insight, leading to wiser maneuvering decisions that help prevent traffic accidents as depicted in Fig. 1. (a). By overcoming the limitations of single-agent perception through the exchange of processed sensing data among multiple agents, challenges like non-line-of-sight (NLOS) occluded blind zones, partial object detection and limited detection range can be significantly mitigated. Typically, cooperative perception relies on the integration of V2X wireless communication, sensor data processing, and fusion modules to create a unified collaboration platform. Beyond intelligent traffic systems, cooperative perception can also be applied in multi-robot automation and multi-UAV (unmanned aerial vehicle) deployment for business applications such as smart logistics warehousing, factory production lines, 3D scene scanning, or panoramic photography.
@@
Meanwhile, in the realm of cooperative perception, existing methods have predominantly focused on fusing single unified modality data (e.g., LiDAR or camera) across multiple agents as depicted in Fig. 1. (b). It was not until the recent breakthroughs with HM-ViT [1] and HEAL [2] that certain forms of heterogeneous fusion became possible as depicted in Fig. 1. (c). However, both HM-ViT and HEAL still restrict themselves by sharing one modality of sensor data per agent only, meaning the full potential of multi-modality remains untapped and invalidated in multi-agent collaboration. Simply forcing fusion of multi-modality sensor data (e.g., LiDAR and camera) per agent on HM-ViT or HEAL could even lead to markedly degraded performance compared to LiDAR-only multi-agent setting, due to the poor depth estimation from the camera modality. Thus, the gap between multi-agent single-modality and full multi-agent multi-modality still exists. Moreover, alongside efforts to close this gap, maintaining fusion compatibility with single-modality agents is crucial, raising the question: how can we fully leverage every possible sensor on every possible agent for cooperative perception? While achieving high perception performance is the primary goal, we also recognize the importance of computational efficiency and communication support-ability for future industrial applications.
@@
In this paper, we propose two novel multi-agent and multi-modality cooperative perception architectures: Paint-To-Puzzle (PTP) and Co-Sketching-Co-Coloring (CoS-CoCo), each emphasizing different performance aspects based on the aforementioned motivations. PTP involves only one single inter-agent fusion step across all agents, with LiDAR modality as a mandatory requirement. This means that each participant performs its own cross-modality fusion before contributing to a unified fusion with all other agents. In contrast, CoS-CoCo uses a two-stage inter-agent fusion process: first fusion of LiDAR feature data among all agents equipped with LiDAR sensors, followed by fusion of camera feature data among all agents with camera sensors. Obviously, CoS-CoCo welcomes any agent with any sensor modality configuration to contribute. The choice between single-stage or two-stage cross-agent fusion affects the trade-off between PTP’s lower communication bandwidth requirements and CoS-CoCo’s broader diversity of sensor-sharing candidates, which leads to differences in perception performance. To enable efficient multi-modality fusion for both architectures, we design Radian-Glue-Attention (RG-Attn). RG-Attn samples the Bird’s Eye View (BEV) feature map constructed from LiDAR sensor data based on the divided radians of the camera’s field-of-view (FOV), enabling linear attention to each corresponding column of the camera feature map. This approach enhances the BEV feature map by incorporating the projected semantic information. Both PTP and CoS-CoCo outperform existing methods in the cooperative perception field by leveraging all available sensor sources as depicted in Fig. 1. (d), achieving dominant performance while significantly reducing training costs compared to most methods and maintaining real-time inference computational efficiency. In general, our contributions are as follows:
@@ II-B Multi-Agent Cooperative Perception
Within the cooperative perception framework, Who2Comm series [6] focused on the priority of messages to transfer, FedBEVT [25] explored federated learning for cooperative perception, while Coopernaut and ICOP [4, 5] demonstrated how cooperative perception could enhance end-to-end (E2E) autonomous driving. The evolution of intermediate fusion methods has spanned concatenation-based, transformer-based, and pyramid-based approaches. Most existing works have focused on the single modality until recent HM-ViT [1] and HEAL [2] introduced multi-modality. However, both HM-ViT and HEAL limited their solution to only one modality of shared data per agent. The existing gap is how much can we improve if fully leverage all possible sensor sources to pursue an oracle-like vision, this critical issue is addressed in this paper.
@@ III METHODOLOGY
We first introduce the LiDAR and camera cross-modality fusion module as depicted in Fig. 2., and then expand on the design of two cooperative perception architectures that implement this fusion module as illustrated in Fig. 3. (a) and (b).
@@ III-A Radian-Glue-Attention (RG-Attn)
Since the generally inaccurate depth estimation from camera data is avoided, the key to successful cross-modality fusion lies in effectively projecting camera feature semantics onto the BEV feature space generated by the LiDAR backbone. Suppose the solid BEV feature map $F^{\text{bev}}_{j}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$ and camera feature $F^{\text{cam}}_{ik}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}}$ are generated from collaborative agents, where $j$ and $i$ denotes the sequence number of agents and $ik$ identify the camera $k$ equipped on agent $i$. Since we do not limit the fusion to occur within a single system, the transformation matrix $T_{i\rightarrow j}$ of agent $i$ to agent $j$ considering the BEV gird size in reality world, along with the sensor $k$ to its boarding agent transform $t_{ik\rightarrow i}$ are utilized for locating the camera sensor’s location $t_{ik\rightarrow j}$ on the BEV feature map of agent $j$ as $T_{i\rightarrow j}\times t_{ik\rightarrow i}$. The rotation part $R_{i\rightarrow j}$ out of transformation matrix $T_{i\rightarrow j}$ and the rotation matrix $R_{ik\rightarrow i}$ of camera $k$ to its mounted agent $i$ to form $R_{ik\rightarrow j}$ as $R_{i\rightarrow j}\cdot R_{ik\rightarrow i}$, will transfer the horizontal FOV angular interval $\left[-\frac{\theta_{\text{FOV}}}{2},\frac{\theta_{\text{FOV}}}{2}\right]$ from camera k on agent i to agent j $\left[\theta^{\text{start}}_{ik\rightarrow j},\theta^{\text{end}}_{ik\rightarrow j}\right]$ as $\left[\text{atan}\left(R_{ik\rightarrow j}\cdot u\left(-\frac{\theta_{\text{FOV}}}{2}\right)\right),\text{atan}\left(R_{ik\rightarrow j}\cdot u\left(\frac{\theta_{\text{FOV}}}{2}\right)\right)\right]$, where $u$ stands for the angle unit vector.
@@
After determining the relative transform and the horizontal FOV range of camera $k$ from agent $i$ on the target BEV map of agent $j$, the most fundamental and intuitive geometric relationship between the camera feature matrix and the intersecting BEV feature map within its perception region is that the camera’s horizontal FOV defines the angular range of its feature matrix projection on the BEV feature map. Therefore, segment the FOV by the width $W_{2}$ of the camera feature matrix $F^{\text{cam}}_{ik}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}}$, generating $W_{2}$ sub-sectors to correspond with each column of the camera feature matrix. Divide each sub-sector along radius by grid number $h$ to create segmented gridded sub-sectors, and define the camera feature fusion maximum range $R$ on the BEV feature map, the correspondence between all the grids of the sector and the overlapping partial grids of BEV feature map could be established by using bilinear sampling, as we just draw a circular gridded sector with fixed radius $R$, coordinates of the sector’s center $(t_{ik\rightarrow j})$, central angle of the sector $\left[\theta^{\text{start}}_{ik\rightarrow j},\theta^{\text{end}}_{ik\rightarrow j}\right]$, angular segment granularity $W_{2}$ and the radial segment granularity $h$ on the target BEV map for later sampling and reverse sampling purpose. The camera feature fusion maximum range $R$ is defined in terms of the granularity of the BEV grid’s side length, which we set to half of the BEV rectangle’s diagonal length to balance distortion levels and fusion range, and we set grid number $h$ to $H_{1}$. All the correspondence parameters form the sampling configuration set ${Set}_{ik\rightarrow j}$ as $(t_{ik\rightarrow j},\ \left[\theta^{\text{start}}_{ik\rightarrow j},\theta^{\text{end}}_{ik\rightarrow j}\right],\ W_{2},\ h,\ R)$.
@@ III-B Paint-To-Puzzle (PTP)
The core concept in the design of PTP is that every agent strives to create the most accurate fused BEV feature map before engaging in cross-agent fusion. As depicted in Fig. 3. (a), agents equipped with both LiDAR and cameras “paint” individual perceived scenarios and then “puzzle” them together. Consequently, the RG-Attn cross-modality fusion module is applied only within single-agent system in PTP, creating $F^{\text{fus-bev}}_{ik\rightarrow i}$ multiple times, depending on the number $n$ of cameras equipped on agent $i$. These semantically enriched features are then concatenated to form $F^{\text{fus-bev}}_{i+\sum_{k=1}^{n}ik}$ as:
@@ III-C Co-Sketching-Co-Coloring (CoS-CoCo)
As depicted in Fig. 3. (b), CoS-CoCo could be clearly divided into two stages of fusion: Co-Sketching, which involves the fusion of LiDAR BEV feature maps among LiDAR-equipped agents to provide the basic skeleton of the perceived surrounding environment, and Co-Coloring, which projects camera features onto the skeleton formed in the first stage. Similar to the BEV feature map fusion in PTP, the Pyramid Fusion module is employed in the Co-Sketching stage to fuse all available BEV feature maps into a robust, unified BEV feature map $F^{\text{bev}}_{pyr}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$ as:
@@ IV-B Settings
Training configurations: Three loss function regarding the classification, regression and direction are used, incorporating foreground map output into loss calculation. We use Adam optimizer, with an initial learning rate set to 0.2 and decrease by 0.1 from epoch 15 to 25 for OPV2V and from epoch 35 to 40 for DAIR-V2X. On an NVIDIA RTX 6000 Ada, the training process with 30 epochs for both OPV2V and 40 epochs for DAIR-V2X in total no matter the model is PTP or CoS-CoCo, costs approximately 6 hours for training on DAIR-V2X if setting batch size to 4. But for the OPV2V case, the number of agents can reach up to 7 and each vehicle’s four-camera setup significantly increase computational demand, resulting in around 36 hours of training time due to GPU memory constraints.
@@ IV-C Quantitative & Visualization Results
As shown in Table I, we compare the evaluation results of our approaches against best performance results of existing methods under identical conditions (i.e., collaborative agents set to 2 for DAIR-V2X and 4 for OPV2V, with a fixed detection range) to ensure a fair comparison. The only distinction is that the compared methods rely solely on LiDAR data from all collaborative agents for their peak performances as reported in the corresponding published papers, CoBEVT and HEAL present performance degradation when incorporating camera data features compulsively by contrasting the AP30 column of Table I and last column of Table II (where both LiDAR and camera modalities are enabled for both agents). In contrast, our approach leverages the additional positive benefits of camera data, which could be reflected by the comparison between HEAL and our proposed methods as we build the cross-modality fusion module based on the LiDAR-only structure of HEAL (i.e., both of our proposed models deliver the same performance as HEAL if restricted to LiDAR-only cross-agent fusion).
@@
Dataset DAIR-V2X (AP30)
Modality LC LC+C LC+L LC+LC
CoBEVT 0.146 0.553 0.589 0.588
HEAL 0.237 0.574 0.692 0.776
PTP (Ours) 0.707 / 0.743 0.862
CoS-CoCo (Ours) 0.705 0.712 0.848 0.854
@@
We further assess our method’s performance across various combinations of agent numbers and modalities on the DAIR-V2X dataset, as presented in Table II, where “+” separates two agents, “L” stands for LiDAR, “C” for camera, and “LC” for agent equipped with both LiDAR and camera. All models are trained only once with full multi-modality multi-agent setting (i,e,. LC+LC), and directly used for all the inferences with different settings. The performances of HEAL and CoBEVT for conducting multi-modality, are achieved by their BEV fusion module to fuse all BEV feature maps from different modalities and agents, decreased by 5.6% and 19.9% in DAIR-V2X AP30 compared to LiDAR-only case, reflecting their negatives of incorporating camera modality.
@@
In addition to the quantitative results, visualizations of the intermediate cross-modality fusion via the RG-Attn module and the final object detection outcomes of our methods versus the previous SOTA method are shown in Fig. 4., respectively.

### v2

@@ Abstract
Cooperative perception offers an optimal solution to overcome the perception limitations of single-agent systems by leveraging Vehicle-to-Everything (V2X) communication for data sharing and fusion across multiple agents. However, most existing approaches focus on single-modality data exchange, limiting the potential of both homogeneous and heterogeneous fusion across agents. This overlooks the opportunity to utilize multi-modality data per agent, restricting the system’s performance. In the automotive industry, manufacturers adopt diverse sensor configurations, resulting in heterogeneous combinations of sensor modalities across agents. To harness the potential of every possible data source for optimal performance, we design a robust LiDAR and camera cross-modality fusion module, Radian-Glue-Attention (RG-Attn), applicable to both intra-agent cross-modality fusion and inter-agent cross-modality fusion scenarios, owing to the convenient coordinate conversion by transformation matrix and the unified sampling/inversion mechanism. We also propose two different architectures, named Paint-To-Puzzle (PTP) and Co-Sketching-Co-Coloring (CoS-CoCo), for conducting cooperative perception. PTP aims for maximum precision performance and achieves smaller data packet size by limiting cross-agent fusion to a single instance, but requiring all participants to be equipped with LiDAR. In contrast, CoS-CoCo supports agents with any configuration—LiDAR-only, camera-only, or LiDAR-camera-both, presenting more generalization ability. Our approach achieves state-of-the-art (SOTA) performance on both real and simulated cooperative perception datasets. The code is now available at GitHub.
@@ I INTRODUCTION
The famous phrase “United we stand, divided we fall” by Aesop aptly captures the essence of multi-agent cooperative perception. Shared and fused perception information can be the crucial stepping stone, akin to an oracle’s insight, leading to wiser maneuvering decisions that help prevent traffic accidents as depicted in Fig. 1 (a). By overcoming the limitations of single-agent perception through the exchange of processed sensing data among multiple agents, challenges like non-line-of-sight (NLOS) occluded blind zones, partial object detection and limited detection range can be significantly mitigated. Typically, cooperative perception relies on the integration of V2X wireless communication, sensor data processing, and fusion modules to create a unified collaboration platform. Beyond intelligent traffic systems, cooperative perception can also be applied in multi-robot automation and multi-UAV (unmanned aerial vehicle) deployment for business applications such as smart logistics warehousing, factory production lines, 3D scene scanning, or panoramic photography.
@@
Meanwhile, in the realm of cooperative perception, existing methods have predominantly focused on fusing single unified modality data (e.g., LiDAR or camera) across multiple agents as depicted in Fig. 1 (b). It was not until the recent breakthroughs with HM-ViT [1] and HEAL [2] that certain forms of heterogeneous fusion became possible as depicted in Fig. 1 (c). However, both HM-ViT and HEAL still restrict themselves by sharing one modality of sensor data per agent only, meaning the full potential of multi-modality remains untapped and invalidated in multi-agent collaboration. Simply forcing fusion of multi-modality sensor data (e.g., LiDAR and camera) per agent on HM-ViT or HEAL could even lead to markedly degraded performance compared to LiDAR-only multi-agent setting, due to the poor depth estimation from the camera modality. Thus, the gap between multi-agent single-modality and full multi-agent multi-modality still exists. Moreover, alongside efforts to close this gap, maintaining fusion compatibility with single-modality agents is crucial, raising the question: how can we fully leverage every possible sensor on every possible agent for cooperative perception? While achieving high perception performance is the primary goal, we also recognize the importance of computational efficiency and communication support-ability for future industrial applications.
@@
In this paper, we propose two novel multi-agent and multi-modality cooperative perception architectures: Paint-To-Puzzle (PTP) and Co-Sketching-Co-Coloring (CoS-CoCo), each emphasizing different performance aspects based on the aforementioned motivations. PTP involves only one single inter-agent fusion step across all agents, with LiDAR modality as a mandatory requirement. This means that each participant performs its own cross-modality fusion before contributing to a unified fusion with all other agents. In contrast, CoS-CoCo uses a two-stage inter-agent fusion process: first fusion of LiDAR feature data among all agents equipped with LiDAR sensors, followed by fusion of camera feature data among all agents with camera sensors. Obviously, CoS-CoCo welcomes any agent with any sensor modality configuration to contribute. The choice between single-stage or two-stage cross-agent fusion affects the trade-off between PTP’s lower communication bandwidth requirements and CoS-CoCo’s broader diversity of sensor-sharing candidates, which leads to differences in perception performance. To enable efficient multi-modality fusion for both architectures, we design Radian-Glue-Attention (RG-Attn). RG-Attn samples the Bird’s Eye View (BEV) feature map constructed from LiDAR sensor data based on the divided radians of the camera’s field-of-view (FOV), enabling linear attention to each corresponding column of the camera feature map. This approach enhances the BEV feature map by incorporating the projected semantic information. Both PTP and CoS-CoCo outperform existing methods in the cooperative perception field by leveraging all available sensor sources as depicted in Fig. 1 (d), achieving dominant performance while significantly reducing training costs compared to most methods and maintaining real-time inference computational efficiency. In general, our contributions are as follows:
@@ II-B Multi-Agent Cooperative Perception
Within the cooperative perception framework, Who2Comm series [6, 30] focused on the priority of messages to transfer, FedBEVT [25] explored federated learning for cooperative perception, while Coopernaut and ICOP [4, 5] demonstrated how cooperative perception could enhance end-to-end (E2E) autonomous driving. For practical real-world deployment, studies such as CoAlign [26] and CBM [27] have demonstrated significant advancements in minimizing relative localization errors, a critical factor in the coordinate transformation across collaborative agents. Additionally, research from the vehicular wireless communication domain [28, 29, 30] has contributed to optimizing the communication protocol stack, aligning it with the upper-layer requirements of cooperative perception. The evolution of intermediate fusion methods has spanned concatenation-based, transformer-based, and pyramid-based approaches. Most existing works have focused on the single modality until recent HM-ViT [1] and HEAL [2] introduced multi-modality. However, both HM-ViT and HEAL limited their solution to only one modality of shared data per agent. The existing gap is how much can we improve if fully leverage all possible sensor sources efficiently to pursue an oracle-like vision, this critical issue is addressed in this paper.
@@ III METHODOLOGY
We first introduce the LiDAR and camera cross-modality fusion module as depicted in Fig. 2, and then expand on the design of two cooperative perception architectures that implement this fusion module as illustrated in Fig. 3 (a) and (b).
@@ III-A Radian-Glue-Attention (RG-Attn)
Since the generally inaccurate depth estimation from camera data is avoided, the key to successful cross-modality fusion lies in effectively projecting camera feature semantics onto the BEV feature space generated by the LiDAR backbone. Suppose the solid BEV feature map $F^{\text{bev}}_{j}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$ and camera feature $F^{\text{cam}}_{ik}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}}$ are generated from collaborative agents, where $j$ and $i$ denotes the sequence number of agents and $ik$ identify the camera $k$ equipped on agent $i$. Since we do not limit the fusion to occur within a single system, the transformation matrix $T_{i\rightarrow j}$ of agent $i$ to agent $j$ considering the BEV gird size in reality world, along with the sensor $k$ to its boarding agent transform $t_{ik\rightarrow i}$ are utilized for locating the camera sensor’s location $t_{ik\rightarrow j}$ on the BEV feature map of agent $j$ as $T_{i\rightarrow j}\cdot t_{ik\rightarrow i}$. The rotation part $R_{i\rightarrow j}$ out of transformation matrix $T_{i\rightarrow j}$ and the rotation matrix $R_{ik\rightarrow i}$ of camera $k$ to its mounted agent $i$ to form $R_{ik\rightarrow j}$ as $R_{i\rightarrow j}\cdot R_{ik\rightarrow i}$, will transfer the horizontal FOV angular interval $\left[-\frac{\theta_{\text{FOV}}}{2},\frac{\theta_{\text{FOV}}}{2}\right]$ from camera k on agent i to agent j $\left[\theta^{\text{start}}_{ik\rightarrow j},\theta^{\text{end}}_{ik\rightarrow j}\right]$ as $\left[\text{atan}\left(R_{ik\rightarrow j}\cdot u\left(-\frac{\theta_{\text{FOV}}}{2}\right)\right),\text{atan}\left(R_{ik\rightarrow j}\cdot u\left(\frac{\theta_{\text{FOV}}}{2}\right)\right)\right]$, where $u$ stands for the angle unit vector.
@@
Once the relative transform and horizontal FOV range of camera $k$ from agent $i$ are determined on the target BEV map of agent $j$, a fundamental geometric relationship emerges. Specifically, the camera’s horizontal FOV determines the angular span of its projected feature matrix on the BEV feature map. To establish the correspondence between the camera feature matrix and the BEV feature map, we segment the FOV into $W_{2}$ sub-sectors, each corresponding to a column in the camera feature matrix $F^{\text{cam}}_{ik}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}}$. Each sub-sector is further divided along the radius into $h$ grid segments, forming a gridded structure. The maximum fusion range $R$ from camera semantics projection on the BEV feature map is defined based on the BEV grid granularity and is set to half of the BEV rectangle’s diagonal length, balancing distortion and fusion range. The number of radial segments $h$ to divide $R$, is set to $H_{1}$ in our setting. Bilinear sampling is employed to establish the correspondence between the sector grids and their overlapping regions on the BEV feature map. This process involves constructing a circular gridded sector centered at $(t_{ik\rightarrow j})$, with an angular span $\left[\theta^{\text{start}}_{ik\rightarrow j},\theta^{\text{end}}_{ik\rightarrow j}\right]$, angular resolution $W_{2}$, radius length $R$ and radial resolution $h$. The resulting sampling configuration set ${Set}_{ik\rightarrow j}$ is defined as: $(t_{ik\rightarrow j},\ \left[\theta^{\text{start}}_{ik\rightarrow j},\theta^{\text{end}}_{ik\rightarrow j}\right],\ W_{2},\ R,\ h)$.
@@ III-B Paint-To-Puzzle (PTP)
The core concept in the design of PTP is that every agent strives to create the most accurate fused BEV feature map before engaging in cross-agent fusion. As depicted in Fig. 3 (a), agents equipped with both LiDAR and cameras “paint” individual perceived scenarios and then “puzzle” them together. Consequently, the RG-Attn cross-modality fusion module is applied only within single-agent system in PTP, creating $F^{\text{fus-bev}}_{ik\rightarrow i}$ multiple times, depending on the number $n$ of cameras equipped on agent $i$. These semantically enriched features are then concatenated to form $F^{\text{fus-bev}}_{i+\sum_{k=1}^{n}ik}$ as:
@@ III-C Co-Sketching-Co-Coloring (CoS-CoCo)
As depicted in Fig. 3 (b), CoS-CoCo could be clearly divided into two stages of fusion: Co-Sketching, which involves the fusion of LiDAR BEV feature maps among LiDAR-equipped agents to provide the basic skeleton of the perceived surrounding environment, and Co-Coloring, which projects camera features onto the skeleton formed in the first stage. Similar to the BEV feature map fusion in PTP, the Pyramid Fusion module is employed in the Co-Sketching stage to fuse all available BEV feature maps into a robust, unified BEV feature map $F^{\text{bev}}_{pyr}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$ as:
@@ IV-B Settings
Training configurations: Three loss function regarding the classification, regression and direction are used, incorporating foreground map output into loss calculation. We use Adam optimizer, with an initial learning rate set to 0.002 and decrease by 0.1 from epoch 15 to 25 for OPV2V and from epoch 35 to 40 for DAIR-V2X. On an NVIDIA RTX 6000 Ada, the training process with 30 epochs for OPV2V and 40 epochs for DAIR-V2X in total no matter the model is PTP or CoS-CoCo, costs approximately 6 hours for training on DAIR-V2X if setting batch size to 4. But for the OPV2V case, where the number of agents can reach up to 5 and each vehicle’s four-camera setup significantly increase computational demand, resulting in around 36 hours of training time due to GPU memory constraints.
@@ IV-C Quantitative & Visualization Results
As shown in Table I, we compare the evaluation results of our approach with the best reported performances of existing methods [1, 2, 3, 8, 31, 32, 33] under identical conditions. Additionally, the performance results of the listed methods have been validated through our comparison experiments. To ensure a fair comparison, the number of collaborative agents is set to 2 (i.e., always two) for DAIR-V2X and 5 (i.e., maximum five) for OPV2V, with a fixed detection range. The key difference is that the compared methods achieve their peak performance using only LiDAR data from all collaborative agents, as reported in the corresponding papers. Notably, CoBEVT and HEAL exhibit performance degradation when incorporating camera data, as observed in the AP30 column of Table I and the LC+LC column under DAIR-V2X (AP30) of Table II (where both LiDAR and camera modalities are enabled for all agents without extending optimization). In contrast, our approach effectively leverages the complementary advantages of camera data. This improvement is particularly evident when comparing HEAL with our proposed methods, as our cross-modality fusion module is built upon LiDAR-only architecture of HEAL. Specifically, when restricted to LiDAR-only cross-agent fusion, both of our proposed models achieve the same performance as HEAL.
@@
Dataset DAIR-V2X (AP30) OPV2V (AP50)
Modality LC LC+C LC+L LC+LC LC LC+C LC+L LC+LC
CoBEVT 0.146 0.553 0.589 0.588 0.472 0.604 0.647 0.643
HEAL 0.237 0.574 0.692 0.776 0.581 0.636 0.733 0.854
PTP 0.707 / 0.743 0.862 0.820 / 0.875 0.955
CoS-CoCo 0.705 0.712 0.848 0.854 0.821 0.837 0.944 0.949
Figure 4: AP50 performance of our methods and key baseline approaches under different maximum numbers of participating agents on OPV2V.
Figure 5: AP70 performance of our methods and key baseline approaches under different maximum numbers of participating agents on OPV2V.
@@
To assess the impact of the number of participating agents in cooperative perception on our proposed methods, we conduct a series of controlled experiments on the OPV2V dataset. As illustrated in Fig. 4 and Fig. 5, these experiments systematically evaluate how performance scales with varying numbers of collaborating agents, providing insights into the effectiveness and adaptability of our approach under different cooperation levels.
We further assess our methods’ performances across various combinations of agent numbers and modalities on both the DAIR-V2X dataset and the OPV2V dataset, as presented in Table II, where “+” separates two agents, “L” stands for LiDAR, “C” for camera, and “LC” for agent equipped with both LiDAR and camera. All models are trained only once with full multi-modality multi-agent setting (i,e,. LC+LC), and directly used for all the inferences with different settings. The performances of HEAL and CoBEVT for conducting multi-modality, are achieved by their BEV fusion module to fuse all BEV feature maps from different modalities and agents, decreased by 5.6% and 19.9% in DAIR-V2X AP30 compared to LiDAR-only case, reflecting their negatives of incorporating camera modality. A similar performance degradation could be also observed on OPV2V, with HEAL and CoBEVT experiencing declines of 6.8% and 27.1% in AP50, as reflected in Fig. 4 and the right half of Table II.
@@
In addition to the quantitative results, visualizations of the intermediate cross-modality fusion via the RG-Attn module and the final object detection outcomes of our methods versus the previous SOTA method are shown in Fig. 6, respectively.